# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 6.1 Overview of the symmetry
see Figure 11, and Figure 12 for the one-loop form factor. We take the dual coordinates to in $D$ dimensions as $X^{M}$, $M\in\{0,1,\ldots,D-1\}$, which will allow us to take the $\mu$ terms into account
@@ 6.2 One-loop level
${\rm d}\mathbb{I}_{3}^{(1)}\equiv{\rm d}X_{a}\,\mathbb{I}_{3}^{(1)}={\rm d}X_{a}\frac{{\rm tr}_{-}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{X_{1a}^{2}X_{2a}^{2}X_{3a}^{2}X_{4a}^{2}X_{\bar{1}a}^{2}}\,,$ (109)
@@
${\rm tr}_{-}(X_{\underline{4}1},X_{12},X_{23},X_{34})={1\over 2}{\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})++2i\varepsilon(X_{\underline{4}1},X_{12},X_{23},X_{34})\,.$ (112)
For the parity-even part, we find that the extra term that is proportional to $q\cdot X_{1a}$ in (110) is canceled by summing cycling the external momenta and shifting the internal coordinate
$\sum_{i=1}^{4}{\rm d}^{D}X_{a}\frac{q\cdot(X_{1}-X_{a}){\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{2X_{ia}^{2}X_{i+1,a}^{2}X_{i+2,a}^{2}X_{i+3,a}^{2}X_{\bar{i}a}^{2}}$
$=$ $\sum_{i=1}^{4}{\rm d}^{D}X_{a}\frac{{\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{4X_{i+1,a}^{2}X_{i+2,a}^{2}X_{i+3,a}^{2}}\left(\frac{1}{X_{ia}^{2}}-\frac{1}{X_{\bar{i}a}^{2}}\right)=0\,.$ (113)
@@ 6.3 Two-loop level
When we take the lightlike limit $q^{2}=0$, the two numerators $N_{1}$ and $N_{4}$ simply vanish, since they are proportional to $q^{2}$. Hence, we only need to consider the remaining topologies with the dual coordinates shown in Figure 13, where $X_{\underline{b}}=X_{b}+q$ is introduced for the non-planar topologies.

### v3

@@ 6.1 Overview of the symmetry
see Figure 11 and Figure 12 for the one-loop form factor. We take the dual coordinates in $D$ dimensions as $X^{M}$, $M\in\{0,1,\ldots,D-1\}$, which will allow us to take the $\mu$ terms into account
@@ 6.2 One-loop level
${\rm d}\mathbb{I}_{3}^{(1)}\equiv{\rm d}^{D}X_{a}\,\mathbb{I}_{3}^{(1)}={\rm d}^{D}X_{a}\frac{{\rm tr}_{-}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{X_{1a}^{2}X_{2a}^{2}X_{3a}^{2}X_{4a}^{2}X_{\bar{1}a}^{2}}\,,$ (111)
@@
${\rm tr}_{-}(X_{\underline{4}1},X_{12},X_{23},X_{34})={1\over 2}{\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})+2i\varepsilon(X_{\underline{4}1},X_{12},X_{23},X_{34})\,.$ (114)
For the parity-even part, we find that the extra term that is proportional to $q\cdot X_{1a}$ in (112) is canceled by summing cycling the external momenta as
$\sum_{i=1}^{4}{\rm d}^{D}X_{a}\frac{q\cdot(X_{i}-X_{a}){\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{2X_{ia}^{2}X_{i+1,a}^{2}X_{i+2,a}^{2}X_{i+3,a}^{2}X_{\bar{i}a}^{2}}$
$=$ $\sum_{i=1}^{4}{\rm d}^{D}X_{a}\frac{{\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{4X_{i+1,a}^{2}X_{i+2,a}^{2}X_{i+3,a}^{2}}\left(\frac{1}{X_{ia}^{2}}-\frac{1}{X_{\bar{i}a}^{2}}\right)$
$=$ ${\rm d}^{D}X_{a}\frac{{\rm tr}(X_{\underline{4}1},X_{12},X_{23},X_{34})\mu_{aa}}{4}\left(\frac{1}{X_{1a}^{2}X_{2a}^{2}X_{3a}^{2}X_{4a}^{2}}-\frac{1}{X_{\bar{1}a}^{2}X_{\bar{2}a}^{2}X_{\bar{3}a}^{2}X_{\bar{4}a}^{2}}\right)$
$=$ $0\,,$ (115)
where $X_{5}\equiv X_{\bar{1}}$, etc. In the third line of the above formula, we shift the internal coordinate $X_{i}$ in the first integrand by $q$, which leads to $X_{ia}^{2}\to X_{\bar{i}a}^{2}$. This shift can be understood as a relabelling of the propagators, which identifies the same Feynman integrands which have the different propagators labels. Since there is no need to perform any integration, we refer to the above operation as a property at the integrand level.
@@ 6.3 Two-loop level
When we take the lightlike limit $q^{2}=0$, the two numerators $N_{1}$ and $N_{4}$ simply vanish, since they are proportional to $q^{2}$. Hence, we only need to consider the remaining topologies with the dual coordinates shown in Figure 13, where $X_{\underline{b}}=X_{b}+q$ is introduced for the non-planar topologies. We comment here that the labels of the dual coordinates are not unique. However, the symmetry property should be true for general assignments of dual coordinates. Different assignments can be understood as a relabelling of the propagators, while the property of DDCI is independent of the particular choice.